# Education-Oriented Graph Retrieval-Augmented Generation for Learning Path Recommendation

### v1

@@ 4.1. KC Textual Analysis and Generation
[…]
Here, we focus on the task of KC analysis optimization via TextGrad. In KC analysis optimization, the goal is to improve the analysis of overall KCs by continuously updating and optimizing the explanatory text for each KC, utilizing LLMs (e.g., gpt-4o) to generate more accurate descriptions. In particular, we often have a computation graph like the following:
(3) $\text{Objective}=\text{LLM}(\text{Question}+\text{Analysis}+\text{Evaluation}\,\,\text{Instruction}),$
where the parameter we optimize is the Analysis, and the loss function is obtained by an evaluation of the analysis.
In this process, the knowledge concept name serves as the input. Through prompt engineering, the LLM is guided to act as an expert in the education domain. It generates an initial explanation for the knowledge concept, which is subsequently evaluated by another LLM. Based on the evaluation results, further refinements are made. This iterative cycle of generating and evaluating explanations continues until an accurate and reliable explanation is produced. For instance, the objective function for the KC analysis is shown in Figure 7.
4.2. Education-Oriented Graph RAG
To automatically generate a knowledge graph that captures the prerequisite and similarity relationships among KCs, we present EDU-Graph RAG, an education-oriented adaptation of Graph RAG (Edge et al., 2024). Graph RAG aims to extract structured data from unstructured text using LLM and build labeled knowledge graphs to support various application scenarios, such as dataset question generation and summary question answering. It provides fine-tuning commands that allow adjustments to its prompt content for specific domains. By inputting customized instructions, Graph RAG is tailored to specialize in analyzing the relationships between KCs in the education field. Moreover, prompt adjustments ensure it can accurately capture prerequisite relationships and similarity relationships among KCs, improving its domain-specific effectiveness. EDU-Graph RAG is tailored specifically for the education domain by fine-tuning its prompt to process the explanatory texts of KCs and extract prerequisite relationships and similarity relationships. The operational process of EDU-Graph RAG is outlined as follows:
Text Chunking. The explanatory texts of knowledge concepts are divided into smaller, manageable chunks suitable for LLM processing. This enhances both recall and precision in analyzing knowledge concept texts.
Extraction of Element Instances. Entities and associated statements are extracted from the parsed explanatory texts. For example, this includes the professional domain of the KC and its prerequisite relationships and similarity relationships with other KCs.
Generation of Element Summaries. Based on the extracted entities and statements, meaningful summaries are created. These summaries may include implied relationships and conceptual overv-
iews that are not explicitly stated in the text.
KC Structure Graph Construction. Finally, using the extracted entities, statements, and summaries, a comprehensive knowledge graph of prerequisite relationships and similarity relationships among KCs is constructed.
This process ensures an accurate and structured representation of prerequisite relationships and similarity relationships, providing the necessary KC structure graph for the learning path recommendation model, which in turn generates personalized learning paths based on learners’ needs and objectives.
5. Learning Path Generation
In this section, we develop a Discrimination Learning-driven Reinforcement Learning (DLRL) framework that alleviates the blocked phenomenon, further improving the performance of learning path recommendations. To recommend an effective learning path, we incorporate a Difficulty Matching Knowledge Tracing (DIMKT) model (Shen et al., 2022) as the environment KT model to dynamically track the evolving knowledge states of learners. Furthermore, we introduce a dual-agent approach that integrates a Prerequisite Agent (P-Agent) and a Similarity Agent (S-Agent), which work collaboratively to monitor the learner’s progress and adaptively select KCs based on either prerequisite or similarity relationships. And a Difficulty Agent (D-Agent), which operates by analyzing the selected KCs and the learner’s mastery level of these KCs. It selects exercises that align with the learner’s current difficulty level while maintaining a consistent and gradual progression in the difficulty of the recommended learning path. This ensures that the learning path remains smooth and appropriately challenging, maximizing the learner’s engagement and effectiveness.
5.1. Knowledge Tracing
The dynamic changes in a learner’s knowledge mastery level during the learning process serve as a reflection of their learning outcomes. By leveraging these insights to devise subsequent recommendation strategies, it is possible to provide a learning path that aligns with the learner’s abilities and progress. Given the maturity and robustness of existing knowledge tracing (KT) methods, such as the PYKT knowledge tracing library (Liu et al., 2022), which allows for the selection of suitable KT approaches to predict learners’ knowledge states, we adopt the DIMKT method for tracking learners’ knowledge mastery level dynamics. DIMKT is particularly advantageous as it can predict the impact of exercise difficulty on learners, thereby enhancing the accuracy of recommending exercises based on their difficulty level.
5.2. Prerequisite Agent
The P-agent is designed to recommend a suitable learning path for learners based on KCs prerequisite relationships. However, due to the vast number of knowledge concepts, this selection process results in a large search space. To reduce this complexity, it is essential to identify a preliminary shortest learning path between an initial node and the learning objective before the P-agent performs its selection. Existing advanced methods, such as adaptive action space learning algorithms, employ the A* algorithm (Duchoň et al., 2014) to generate the shortest paths while dynamically determining candidate action spaces. Nonetheless, these methods often fail to identify an effective initial node, which may result in the selection of unsuitable knowledge concepts at the outset and ultimately lead to an ineffective learning path.
To address this limitation, we introduce an initial node identification algorithm to ensure the selection of the most appropriate initial node. Specifically, our approach begins from the target knowledge concept node and iteratively traces back along the prerequisite relationship. During this backward traversal, we assess whether the prerequisite knowledge concepts of the current node have already been mastered by the learner. If these prerequisite concepts are mastered, the backward traversal terminates. Otherwise, we continue tracing back to the node with the longest chain of prerequisite knowledge concepts. By incorporating this method, the most suitable initial node can be determined before dynamically identifying the candidate action space, effectively optimizing the overall learning path selection process.
5.2.1. State encoder.
At each step, the predicted knowledge state is based on the learner’s learning objectives and historical learning records. Thus, at the end of the $t-1$ step, the state of the P-agent includes the learner’s knowledge state $\mathbf{h}_{t-1}$ and learning objectives $\mathbf{G}$. Specifically, we use a one-hot encoding scheme to represent the learning objective $\mathbf{G}=\{0,1\}^{M}$, where $M$ is the total number of knowledge concepts. The index corresponding to the learning objective is set to $1$, while all other indices are set to 00. At step $t$, the state of the P-agent $\mathbf{s}_{t}$ is represented as follows:
(4) $\mathbf{s}_{t}=\mathbf{h}_{t-1}\oplus\mathbf{G}.$
5.2.2. Policy.
We adopt Proximal Policy Optimization (PPO) (Schulman et al., 2017) as the model for the prerequisite relation agent, as it is one of the most performant reinforcement learning models currently available. PPO is particularly effective at identifying the optimal sequence of knowledge concepts to achieve the learning goal. Specifically, PPO consists of two primary components: a policy network (Actor) $\pi(\mathbf{G}\left|\mathbf{s}_{t};\theta\right.)$ and a value network (Critic) $V(\mathbf{s}_{t};\phi)$, where $\theta$ and $\phi$ represent the respective network parameters.
@@ 5.2. Prerequisite Agent
The goal of the policy network is to generate a probability distribution over the knowledge concepts in the dataset. The state of the P-agent is fed into the policy network, which outputs a corresponding probability distribution over the knowledge concepts.
(5) $\mathbf{G}\sim\pi(\mathbf{G}\left|\mathbf{s}_{t};\theta\right.)=\text{Softmax}(\text{FC}(\mathbf{s}_{t})),$
@@
where $FC$ is the fully connected layer. The value network takes the state of the relation agent as input and evaluates the reward associated with that state.
(6) $V\left(\mathbf{s}_{t};\phi\right)=\text{FC}\left(\mathbf{s}_{t}\right).$
@@
where $Re_{t}$ represents the reward at step $t$, the computation of which will be detailed later, and $\gamma$ denotes the discount factor. Based on existing reinforcement learning optimization algorithms for sequential decision-making, we optimized the policy network using PPO-clip (Chen et al., 2023).
(8) $\hat{A}_{t}=-V\left(\mathbf{s}_{t};\phi\right)+Re_{t}+\cdots+\gamma^{T-t+1}Re_{T-1}+\gamma^{T-t}V\left(\mathbf{s}_{T};\phi\right),$
(9) $L(\theta)=\hat{\mathbb{E}}_{t}\left[\text{min}\left(r_{t}\left(\theta\right)\hat{A}_{t},clip\left(r_{t}\left(\theta\right),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right],$
where $r_{t}(\theta)$ represents the probability ratio $\frac{\pi(\left.G\right|s_{t};\theta)}{\pi\left(\left.G\right|s_{t};\theta_{old}\right)}$, $\epsilon=0.2$ is a hyperparameter, and $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}$ modifies the surrogate objective by clipping the probability ratio, which removes the incentive for moving $r_{t}$ outside of the interval $[1-\epsilon,1+\epsilon]$.
5.2.3. Reward
To maximize the improvement in learners’ mastery of the learning objectives, we calculate the corresponding reward value after selecting knowledge concepts at each step, as described in Section 3.1.
(10) $Re_{t}=\begin{cases}E_{p},\,\,\text{if t is the last learning stage}\\ 0,\,\,\,\,\text{otherwise.}\\ \end{cases}$
5.3. Similarity Agent
While the prerequisite relationships between knowledge concepts provide a basis for recommending a learning path, they often exhibit limited efficacy in significantly enhancing a learner’s understanding of the current knowledge concept. To address this issue, we introduced a similarity agent (S-agent). The S-agent aims to identify knowledge concepts that are closely related to the target knowledge concept. These similar knowledge concepts can effectively support the learner’s comprehension by reinforcing their grasp of the target knowledge concept. By leveraging the similarity relationship, the S-agent ensures that the recommended learning path not only adheres to prerequisite relationships but also optimally promotes the learner’s mastery of the target knowledge concepts.
At step $t$, the P-agent evaluates the improvement in the learner’s mastery of prerequisite knowledge concept $c_{t-1}$ when selecting the next knowledge concept $c_{t}$ to recommend. The evaluation is based on the following:
(11) $p_{t}=h_{t-1}^{c_{t}}-h_{t-2}^{c_{t}},$
where $h_{t-1}^{c_{t}}$ represents the learner’s knowledge mastery level associated with the knowledge concept $c_{t}$ learned at step $t-1$, and $h_{t-2}^{c_{t}}$ represents the knowledge mastery level at step $t-2$.
If the improvement $p_{t}$ is below a predefined threshold $\tau$, the S-agent is activated. The S-agent identifies a set of similar knowledge concepts that are related to the current knowledge concept $c_{t}$. From this set, the S-agent selects a sub-path that can effectively enhance the learner’s mastery of the current knowledge concept. This sub-path is then appended to the learning path generated by the P-agent. Once the S-agent completes its selecting process, the P-agent resumes its operation, selecting subsequent knowledge concepts based on their prerequisite relationship.
Table 1. Performance comparison for learning path recommendation methods. Existing state-of-the-art results are underlined and the best results are bold. Our DLELP is compared with the SOTA DLPR and * indicates a $p-value<0.05$ in the t-test.
Dataset Steps KNN GRU4Rec RL-Tutor Actor-Critic CSEAL SRC GEHRL DLPR DLELP
@@ 5.3. Similarity Agent
5.4. Difficulty Agent
After the P-Agent and S-Agent select the knowledge concepts, we incorporate a difficulty agent to match the selected concepts with exercises that align with the learner’s knowledge mastery level. This approach prevents the recommended exercises from being too difficult or too easy, which could lead to a loss of interest during the learning process (Huang et al., 2017; Minn et al., 2018; Zhang et al., 2022). Each knowledge concept may correspond to one or more exercises with varying levels of difficulty. For knowledge concepts where the learner’s mastery is low, easier exercises are required. Conversely, for concepts where mastery is high, more challenging exercises are necessary. Therefore, we need to find the exercise $e_{m}$ whose difficulty is closest to the learner’s knowledge state.
@@ 5.4. Difficulty Agent
[…]
@@ 6.1. Datasets
We conduct our experiments using three publicly available educational datasets to validate the effectiveness of our model: Junyi  [^1] https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=1198, MOOCCubeX  [^2] https://github.com/THU-KEG/MOOCCubeX - Computer, and ASSISTments2009  [^3] https://sites.google.com/site/assistmentsdata/home/2009-2010-assistment-data (ASSIST09). All three educational datasets include learner activity data, which comprises learning records as well as the corresponding knowledge concept names. Please note that the Junyi dataset provides a comprehensive KC structure graph, illustrating the prerequisite relationships among knowledge concepts. In contrast, the MOOCCubex dataset contains an incomplete KC structure graph, while the ASSIST09 dataset lacks a KC structure graph altogether. Furthermore, none of the datasets include information on the similarity relationships between KCs. However, all three datasets contain fields with knowledge concept names, allowing us to construct KC structure graphs for both prerequisite and similarity relationships using our method. The statistical details of the datasets are shown in Table 2.
@@
Dataset Junyi MOOCCubeX ASSIST09
@@ 6.3. Implementation Detail
We implemented our proposed learning path recommendation framework using PyTorch (Paszke et al., 2019) and the simulator code from previous works(Li et al., 2023; Zhang et al., 2024). In the process of generating KC structures graph, we adjusted the value of the chunk based on the number of KCs in the dataset. The chunk values used for the Junyi, Assist09, and MOOCCubex datasets were $800$, $600$, and $400$, respectively. In the process of recommending learning path, the first 60% of exercise logs in the dataset were used to train the environment KT, while the remaining exercise logs were used to train and test our model. To more accurately evaluate the model’s performance, we trained a DKT (Piech et al., 2015) on the entire dataset to predict learners’ mastery of target knowledge concepts in Eq. (1). Further, similar to previous methods (Zhang et al., 2024; Li et al., 2023; Liu et al., 2019), we considered a learner’s mastery of a KC to be greater than $0.5$, as this implies a probability greater than $0.5$ of correctly answering an exericise related to that KC. Through testing, we observed that when using our DKT (Piech et al., 2015) for evaluation, the average improvement in learners’ knowledge state per exercise was $0.001$. Thus, in section 5.3, we set the threshold $\tau$ to 0.001. We used the Adam optimizer (Diederik and Ba, 2015) for both the policy network and the value network, with a learning rate of 0.001. Our code is available at https://anonymous.4open.science/r/DLELP.
@@ 6.4. Overall Performance
Table 1 presents the overall performance of all methods across the three datasets, revealing several key insights. In comparison with all baseline models, our method outperforms the others on all three datasets. Specifically, among all learning path recommendation methods, our approach surpasses KC correlation-based methods (Chen et al., 2023), as well as KC prerequisite relation-based methods that consider knowledge concept structures and changes in learners’ knowledge states during sequential decision-making, such as CSEAL (Liu et al., 2019), GEHRL (Li et al., 2023), and DLPR (Zhang et al., 2024). These advanced prerequisite relation-based methods are outperformed by our approach, demonstrating the progress our method makes compared to previous methods. Furthermore, in each dataset, performance generally fluctuates with the variation in recommended steps. This indicates that, aside from the fact that recommending different exercises can have varying effects on learners, the number of recommended exercises also plays a role. When step=20, the recommended learning paths achieved optimal effects across all three datasets. This indicates that the learning path of this length is relatively suitable for the overall learners, enabling them to better grasp the target knowledge concept during the learning process.
Table 3. Ablation Experiment: ”w/o S” represents the model without similarity agent, and ”w S” represents the overall model.
Junyi MOOCCubeX ASSIST09
@@ 6.5. Ablation Study
To validate the effectiveness of the innovative components of our method, we conducted an ablation study. Specifically, we removed the similarity agent module from our method to assess its contribution to the overall performance. As shown in Table LABEL:tab:Ablation_Experiment, when the similarity agent is removed from our method, there is a noticeable performance drop across all three datasets. This highlights the necessity and value of considering the similarity relationships between knowledge concepts in learning path recommendations.
Figure 4. Comparison between the original graph and the graph generated by DLELP. KC denotes the number of knowledge concepts, while Relationship represents the number of prerequisite relationships among these KCs.
Figure 5. Performance comparison of the original graph and the DLELP-generated graph.

### v2

@@ 4.2.2. Prerequisite Agent
The goal of the policy network is to generate a probability distribution over the knowledge concepts in the dataset. The state of the P-agent is fed into the policy network, which outputs a corresponding probability distribution over the KCs.
(8) $\mathbf{G}\sim\pi(\mathbf{G}\left|\mathbf{s}_{t};\theta\right.)=\mathrm{Softmax}(\mathrm{Linear}(\mathbf{s}_{t})),$
@@
where $\mathrm{Linear}$ is the fully connected layer. The value network takes the state of the relation agent as input and evaluates the reward associated with that state.
(9) $V\left(\mathbf{s}_{t};\phi\right)=\mathrm{Linear}\left(\mathbf{s}_{t}\right).$
@@
where $Re_{t}$ represents the reward at step $t$, the computation of which will be detailed later, and $\gamma$ denotes the discount factor. Based on existing reinforcement learning optimization algorithms, we optimized the policy network using PPO-clip (Chen et al., 2023).
(11) $\hat{A}_{t}=-V\left(\mathbf{s}_{t};\phi\right)+Re_{t}+\cdots+\gamma^{T-t}V\left(\mathbf{s}_{T};\phi\right),$
(12) $L(\theta)=\hat{\mathbb{E}}_{t}\left[\mathrm{min}\left(r_{t}\left(\theta\right)\hat{A}_{t},\mathrm{clip}\left(r_{t}\left(\theta\right),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right],$
where $r_{t}(\theta)$ represents the probability ratio $\frac{\pi(\left.G\right|s_{t};\theta)}{\pi\left(\left.G\right|s_{t};\theta_{old}\right)}$, $\epsilon=0.2$ is a hyperparameter, and $\mathrm{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}$ modifies the surrogate objective by clipping the probability ratio, which removes the incentive for moving $r_{t}$ outside of the interval $1-\epsilon,1+\epsilon]$.
c) Reward: To maximize the improvement in learners’ mastery of the learning objectives, we calculate the corresponding reward value after selecting knowledge concepts at each step.
[…]
Table 1. Performance comparison for learning path recommendation methods. Existing state-of-the-art results are underlined and the best results are bold. Our KnowLP is compared with the SOTA DLPR and * indicates a $p-\text{value}<0.05$ in the t-test.
Dataset Steps KNN GRU4Rec RL-Tutor Actor-Critic CSEAL SRC GEHRL DLPR KnowLP
@@
4.2.3. Similarity Agent
While the prerequisite relationships between knowledge concepts provide a basis for recommending a learning path, a learner’s failure to master a key concept may hinder the effective acquisition of subsequent knowledge. To address this issue, we introduced a similarity agent (S-agent). The S-agent identifies a set of similar knowledge concepts that are related to the current knowledge concept $c_{t}$. From this set, the S-agent selects a sub-path that can effectively enhance the learner’s mastery of the current knowledge concept. This sub-path is then appended to the learning path generated by the P-agent. Please note that, except for its use of a similarity-based knowledge concept structure graph as input, the S-agent operates identically to the P-agent in terms of architecture.
4.2.4. Difficulty Agent
To prevent the recommended exercises from being too difficult or too easy, which could lead to a loss of interest during the learning process (Minn et al., 2018; Zhang et al., 2022). We need to find the exercise $e_{m}$ whose difficulty is closest to the learner’s knowledge state.
@@ 4.2.4. Difficulty Agent
where $m$ represents the exercise ID that best matches the difficulty for the learner, the symbol $\left|\cdot\right|$ represents the absolute value, $\text{difficult}(e^{c_{i}})$ represents the difficulty of the exercise $e^{c_{i}}$ corresponding to the KC $c_{i}$, and $h^{c_{i}}$ indicates the learner’s mastery level of the KC $c_{i}$.
4.2.5. Learning Path Construction Mechanism
In this section, we detail the strategy for selecting initial nodes as well as the communication mechanism among agents.
a) Node Initialization: Due to the vast number of knowledge concepts, this selection process results in a large search space. To reduce this complexity, it is essential to identify a preliminary shortest learning path between an initial node and the learning objective before the P-agent performs its selection. Existing advanced methods, such as adaptive action space learning algorithms, employ the A* algorithm (Duchoň et al., 2014) to generate the shortest paths while dynamically determining candidate action spaces. Nonetheless, these methods often fail to identify an effective initial node, which may result in the selection of unsuitable knowledge concepts at the outset and ultimately lead to an ineffective learning path.
To address this limitation, we introduce an initial node identification algorithm to ensure the selection of the most appropriate initial node. Specifically, our approach begins from the target knowledge concept node and iteratively traces back along the prerequisite relationship. During this backward traversal, we assess whether the prerequisite knowledge concepts of the current node have already been mastered by the learner. If these prerequisite concepts are mastered, the backward traversal terminates. Otherwise, we continue tracing back to the node with the longest chain of prerequisite knowledge concepts. By incorporating this method, the most suitable initial node can be determined before dynamically identifying the candidate action space, effectively optimizing the overall learning path generation.
b) Agent Switching: At step $t$, the P-agent evaluates the improvement in the learner’s mastery of prerequisite KC $c_{t-1}$ when selecting the next knowledge concept $c_{t}$ to recommend. The evaluation is based on the following:
(15) $p_{t}=h_{t-1}^{c_{t}}-h_{t-2}^{c_{t}},$
where $h_{t-1}^{c_{t}}$ represents the learner’s knowledge mastery level associated with the knowledge concept $c_{t}$ learned at step t1t-1, and $h_{t-2}^{c_{t}}$ represents the knowledge mastery level at step t2t-2. If the improvement $p_{t}$ is below a predefined threshold $\tau$, the S-agent is activated. Once the S-agent completes its selecting process, the P-agent resumes its operation, selecting subsequent KCs based on their prerequisite relationship. After the P-Agent and S-Agent select the knowledge concepts, we incorporate a difficulty agent to match the selected concepts with exercises that align with the learner’s knowledge mastery level.
@@ 5.1. Datasets
We conduct our experiments using three publicly available educational datasets to validate the effectiveness of our model: Junyi  [^1] https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=1198, MOOCCubeX  [^2] https://github.com/THU-KEG/MOOCCubeX - Computer (MCX), and ASSISTments2009  [^3] https://sites.google.com/site/assistmentsdata/home/2009-2010-assistment-data (ASS09). All three educational datasets include learner activity data, which comprises learning records as well as the corresponding knowledge concept names. Please note that the Junyi dataset provides a comprehensive KC structure graph, illustrating the prerequisite relationships among knowledge concepts. In contrast, the MOOCCubex dataset contains an incomplete KC structure graph, while the ASS09 dataset lacks a KC structure graph altogether. Furthermore, none of the datasets include information on the similarity relationships between KCs. However, all three datasets contain fields with knowledge concept names, allowing us to construct KC structure graphs for both prerequisite and similarity relationships using our method. The statistical details of the datasets are shown in Table 2.
@@
Dataset Junyi MOOCCubeX ASS09
@@ 5.3. Implementation Detail
We implemented our proposed learning path recommendation framework using PyTorch (Paszke et al., 2019) and the simulator code from previous works(Li et al., 2023; Zhang et al., 2024). In the process of generating KC structures graph, we adjusted the value of the chunk based on the number of KCs in the dataset. The chunk values used for the Junyi, ASS09, and MOOCCubex datasets were $800$, $600$, and $400$, respectively. In the process of recommending learning path, the first 60% of exercise logs in the dataset were used to train the environment KT, while the remaining exercise logs were used to train and test our model. To more accurately evaluate the model’s performance, we trained a DKT (Piech et al., 2015) on the entire dataset to predict learners’ mastery of target knowledge concepts in Eq. (1). Further, similar to previous methods (Zhang et al., 2024; Li et al., 2023; Liu et al., 2019), we considered a learner’s mastery of a KC to be greater than $0.5$, as this implies a probability greater than $0.5$ of correctly answering an exericise related to that KC. Through testing, we observed that when using our DKT (Piech et al., 2015) for evaluation, the average improvement in learners’ knowledge state per exercise was $0.001$. Thus, in section 4.2.3, we set the threshold $\tau$ to 0.001. We used the Adam optimizer (Diederik and Ba, 2015) for both the policy network and the value network, with a learning rate of 0.001.
@@ 5.4. Overall Performance
Table 1 presents the overall performance of all methods across the three datasets, revealing several key insights. In comparison with all baseline models, our method outperforms the others on all three datasets. This demonstrates that leveraging LLM-generated knowledge concept graphs, combined with the integration of similarity relationships, contributes significantly to improving recommendation performance. Specifically, EDU-GraphRAG effectively mitigates the hallucination problem associated with knowledge concept structure graphs generated by LLMs, which then serve as reliable guidance for learning path recommendation. Moreover, the incorporation of similarity relationships facilitates smoother learning progression, thereby enhancing the overall performance.
Furthermore, most baseline methods exhibit a decline in performance or fail to achieve their best results when the recommendation step reaches 20. In contrast, KnowLP consistently achieves optimal performance at step=20 across all three datasets. This further demonstrates that the similarity-aware design effectively smooths the learning process.
Table 3. Ablation Experiment: “w/o S” represents the model without similarity agent, and “w S” represents the overall model.
Junyi MOOCCubeX ASS09
@@ 5.5. Ablation Study
To validate the effectiveness of the innovative components of our method, we conducted an ablation study. Specifically, we removed the similarity agent module from our method to assess its contribution to the overall performance. As shown in Table LABEL:tab:Ablation_Experiment, when the similarity agent is removed from our method, there is a noticeable performance drop across all three datasets. This highlights the necessity and value of considering the similarity relationships between knowledge concepts in learning path recommendations. Moreover, with the similarity agent included, our model achieves a substantial performance improvement at step=20. This further confirms that similarity relationships play a critical role in unlocking the performance potential of longer learning paths.
Figure 3. Comparison between the original graph and the graph generated by knowLP. KC denotes the number of knowledge concepts, while Relationship represents the number of prerequisite relationships among these KCs.
Figure 4. Performance comparison of the original graph and the knowLP-generated graph.